On average, an insurance company receives 6 claims between 14:00 and 16:00 on a particular day. What is the probability that the company receives exactly 17 claims between 08:00 and 16:00?

1
Expert's answer
2022-05-24T14:07:40-0400

Let "X=" the number of claims received: "X\sim Po(\lambda t)."

Given "\lambda=6,t=\dfrac{16-8}{16-14}=4"


"P(X=7)=\dfrac{e^{-6\cdot 4}(6\cdot 4)^{17}}{{17}!}\approx0.030864"